Skip to main contentSkip to solution

A random variable X has the following probability distribution:

X0123456
P(X)z2z2z3zz2z^22z22z^27z2+z7z^2+z

Which of the following are TRUE?

(A) P(X < 3) = 0.8

(B) P(X > 4) = 0.2

(C) The value of z is 0.1

(D) P(3≤X<5)=0.31P(3\le X<5) = 0.31

Choose the correct answer from the options given below:

Solution

✅ Correct Option: 4

Sum of probabilities: 10z2+9z=1⇒(10z−1)(z+1)=0⇒z=0.110z^2+9z=1\Rightarrow(10z-1)(z+1)=0\Rightarrow z=0.1, so (C) is true.

P(X<3)=5z=0.5P(X<3)=5z=0.5, so (A) fails.

P(X>4)=2z2+7z2+z=9z2+z=0.19P(X>4)=2z^2+7z^2+z=9z^2+z=0.19, so (B) fails.

P(3≤X<5)=3z+z2=0.30+0.01=0.31P(3\le X<5)=3z+z^2=0.30+0.01=0.31, so (D) is true.

Keyboard Shortcuts

  • Left arrow: Previous question
  • Right arrow: Next question
  • S key: Jump to solution
  • Q key: Jump to question